# Efficient Adjusted Joint Significance Test and Sobel-Type

### v7

@@ Abstract
Mediation analysis is an important statistical tool in many research fields. Its aim is to investigate the mechanism along the causal pathway between an exposure and an outcome. The joint significance test is widely utilized as a prominent statistical approach for examining mediation effects in practical applications. Nevertheless, the limitation of this mediation testing method stems from its conservative Type I error, which reduces its statistical power and imposes certain constraints on its popularity and utility. The joint significance tests for large-scale mediation hypotheses in genome-wide epigenetic research have been extensively investigated, whereas these methods are not applicable to single or small-scale mediation hypotheses. The proposed solution to address this gap is the adjusted joint significance test for one mediator, a novel data-adjusted test for mediation effect that exhibits significant advancements compared to traditional joint significance test. The proposed method is designed to be user-friendly, eliminating the need for complicated procedures. We have derived explicit expressions for size and power, ensuring the theoretical validity of our approach. Furthermore, we extend the adjusted joint significance tests for small-scale mediation hypotheses with family-wise error rate (FWER) control. Additionally, a novel adjusted Sobel-type approach is proposed for the estimation of confidence intervals for the mediation effects, demonstrating significant advancements over conventional Sobel’s confidence intervals in terms of achieving desirable coverage probabilities. Our mediation testing and confidence intervals procedure is evaluated through comprehensive simulations, and compared with numerous existing approaches. Finally, we illustrate the usefulness of our method by analysing three real-world datasets with continuous, binary and time-to-event outcomes, respectively. The method we propose provides a publicly accessible and user-friendly R package, called $\tt AdapMed$, which can be found at https://github.com/zhxmath/AdapMed.
Keywords: Adjusted Sobel’s test; Confidence intervals; Multiple mediators; Small-scale mediation hypotheses; Survival analysis.
@@ 1 Introduction
Mediation analysis plays an important role in understanding the causal mechanism that an independent variable $X$ affects a dependent variable $Y$ through an intermediate variable (mediator) $M$. The application of mediation analysis is widely popular in many fields, such as psychology, economics, epidemiology, medicine, sociology, behavioral science, and many others. From the view point of methodological development, Baron and Kenny (1986) has laid a foundation for the development of mediation analysis. Afterwards, there have been a great deal of papers on this topic. Just to name a few: Mackinnon et al. (2004) constructed the confidence limits for indirect effect by resampling methods. Wang and Zhang (2011) and Zhang and Wang (2013) introduced the estimating and testing methods for mediation effects with censored data and missing data, respectively. Shen et al. (2014) proposed an inference technique for quantile mediation effects. Zhang et al. (2016) introduced a joint significance test method for high-dimensional mediation effects by sure independent screening and minimax concave penalty techniques. VanderWeele and Tchetgen (2017) considered causal mediation analysis with time-varying exposures and mediators. Sun et al. (2021) proposed a Bayesian modeling approach for mediation analysis. Zhou (2022) introduced a semiparametric estimation method for mediation analysis with multiple causally ordered mediators. For more results about mediation analysis, we refer to three reviewing papers by MacKinnon et al. (2007), Preacher (2015) and Zhang et al. (2022).
@@
In the area of mediation analysis, the joint significance test is an important statistical method for investigating the causal mechanism of mediation effects MacKinnon (2008). However, the main shortcoming of this method is due to the conservative type I error of mediation testing MacKinnon et al. (2002), which largely prevents its popularity for practical users. Recently, He et al. (2023) developed a novel adaptive bootstrap testing framework that can be applied to the joint significance test. Furthermore, the confidence interval for the mediation effect is a crucial aspect in mediation analysis, which is highly valuable in comprehending the mediation mechanism. The literature on mediation analysis lacks a substantial number of studies focusing on confidence interval estimation, where the Sobel-type (or normality-based) method is consistently employed to construct these intervals in practical applications. The performance of Sobel-type confidence interval, however, falls short when the mediation effect is zero. To improve the performances of joint significance test and Sobel-type confidence interval, we propose two novel data-adjusted statistical mediation analysis methods with theoretical verification. Specifically, the main advantages of our proposed method are as follows: First, the adjusted joint significance test and adjusted sobel-type confidence interval are two flexible and data-driven methods. Particularly, it is very convenient to implement the two proposed methods from the view of practical application. i.e., our method is user-friendly without involving complicated procedures. Second, compared to the conventional joint significance test, our test method has significantly improved in terms of size and power. The enhanced powers are especially obvious for those relatively weak mediation effects, which are difficult to be recognized as significant mediators by traditional methods. Third, the explicit formulation of the coverage probability for the adjusted Sobel-type confidence interval is provided to ensure the validity of our procedure.
@@
The remainder of this paper is organized as follows: In Section 2, we review some details about the traditional joint significance test for mediation effects. Then we propose a novel data-adjusted joint significance test for one mediator, together with the explicit expressions of size and power. Meanwhile, an adjusted Sobel test is also introduced, which shows a significant improvement compared to traditional Sobel’s method. In Section 3, we implement the adjusted joint significance test towards small-scale multiple testing with FWER control. Section 4 introduces a data-driven approach to construct a confidence interval for the mediation effect. Section 5 presents some simulation studies to assess the performance of our method. In Section 6, we perform mediation analysis for three real-world datasets with our proposed method. Some concluding remarks are provided in Section 7. All proof details of theorems are presented in the Supplementary Material. Finally, our method offers a publicly available and user-friendly R package, called $\tt AdapMed$, which can be accessed at https://github.com/zhxmath/AdapMed.
@@ 2 Adjusted Joint Significance Test
To begin with, we review some basic notations in the framework of mediation analysis. Let $X$ be an exposure, $M$ be the mediator and $Y$ be the outcome (see Figure 1). As described by MacKinnon (2008), the aim of mediation analysis is focused on investigating the causal mechanism along the pathway $X\rightarrow M\rightarrow Y$. Generally speaking, the causal effect $X\rightarrow M$ is parameterized by $\alpha$ (after adjusting for confounders), and the causal effect $M\rightarrow Y$ is parameterized by $\beta$ (after adjusting for exposure and confounders). In this case, the mediating effect of $M$ is described by $\alpha\beta$, which is commonly referred to as the “product-coefficient” approach MacKinnon et al. (2002). To evaluate whether $M$ plays an intermediary role between $X$ and $Y$, it is customary to conduct hypothesis testing at a significance level of $\delta$ as follows:
@@
If the null hypothesis is rejected, we could regard $M$ as a significant mediator along the pathway $X\rightarrow M\rightarrow Y$. It is worth to pointing out that the null hypothesis $H_{0}$ is composite. To be specific, $H_{0}$ can be equivalently decomposed into the union of three disjoint component null hypotheses $H_{0}=H_{00}\cup H_{01}\cup H_{10}$, where
@@
Let $T_{\alpha}$ and $T_{\beta}$ be the statistics for testing $\alpha=0$ and $\beta=0$, respectively. Under the null hypothesis, as the sample size $n$ tends towards infinity, it can be observed that
$T_{\alpha}=\frac{\hat{\alpha}}{\hat{\sigma}_{\alpha}}\sim N(0,1)~{}{\rm and}~{}T_{\beta}=\frac{\hat{\beta}}{\hat{\sigma}_{\beta}}\sim N(0,1),$ (2.2)
where $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively. $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively. The corresponding p-values for $T_{\alpha}$ and $T_{\beta}$ are
@@
where $T_{\alpha}$ and $T_{\beta}$ are defined in (2.2). In the field of mediation analysis, one of the most popular method for (2.1) is the joint significance (JS) test (also known as the MaxP test). The idea of JS test is that we can reject $H_{0}$ if both $\alpha=0$ and $\beta=0$ are simultaneously rejected. i.e., the JS test statistic is defined as
@@
where $P_{\alpha}$ and $P_{\beta}$ are given in (2.3) and (2.4), respectively. Due to the convenience of JS test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the JS test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). In fact, this is a long-existed and unresolved problem for the Sobel test. To bridge this gap, we propose a novel data-adjusted method for improving the statistical efficiency of the Sobel test, especially focusing on the conservative issue in the case of $\alpha=\beta=0$.
Under $H_{0}$, the traditional JS test regards $P_{JS}$ as a uniform random variable over $(0,1)$. i.e., $P_{JS}\sim U(0,1)$. However, the actual distribution of $P_{JS}$ is not $U(0,1)$ under the component null hypotheses $H_{00}$, which is the reason of the conservative performance of traditional JS test. To provide more insights on this issue, suppose that $Z_{1}$ and $Z_{2}$ are two independent random variables following U(0,1) distribution. Let ${Z}=\max({Z_{1}},{Z_{2}})$ and its density function is denoted as $f_{{Z}}(z)$. The distribution function of $Z$ is
@@
i.e., under $H_{00}$ the squared maximum of two p-values ($P^{2}_{JS}$) has an uniform distribution over (0, 1). This finding sheds light on the solution to the conservative issue of conventional JS test.
@@
Following He et al. (2023), we try to reasonably distinguish $H_{00}$ from $H_{01}$ and $H_{10}$ using a certain criterion. Under $H_{00}$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{00})=1$, where $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as the sample size $n\rightarrow\infty$. However, $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{10})=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{01})=0$. In view of these points, we are likely to regard the null hypothesis $H_{0}$ as $H_{00}$ instead of $H_{01}$ or $H_{10}$ when $\max\{|T_{\alpha}|,|T_{\beta}|\}<\lambda_{n}$. Note that He et al. (2023) employed a similar approach in constructing bootstrap-based testing statistics for evaluating mediation effects, whereas our focus lies in the development of novel testing statistics without the need for re-sampling procedures. We provide a numerical example to illustrate the above-mentioned spirit. Specifically, we generate a series of random samples from the linear mediation models: $Y=0.5+0.5X+\beta M+\epsilon$ and $M=0.5+\alpha X+e$, where $X$, $\epsilon$ and $e$ follow from $N(0,1)$. The resulting $T_{\alpha}$ and $T_{\beta}$ are obtained by the ordinary least square method. In Figure 2, we present the scatter plots of $T_{max}=\max\{|T_{\alpha}|,|T_{\beta}|\}$ and $\lambda_{n}=\sqrt{n}/\log(n)$ with 1000 repetitions, where the sample size is $n=1000$, the parameters $(\alpha,\beta)$ are chosen as $(\alpha,\beta)=(0,0)$, and (0.25, 0), respectively. The observation from Figure 2 suggests that, under the assumption of $\alpha=\beta=0$, there is a high probability that $T_{max}$ is significantly smaller than $\lambda_{n}$ as $n\rightarrow\infty$. However, $T_{max}$ is asymptotically much larger than $\lambda_{n}$ when $\alpha\neq 0$, and the same conclusion also applies to $\beta\neq 0$.
Motivated by the above findings, we propose a novel adjusted joint significance (AJS) test procedure for (2.1), where the p-value is defined as
@@
Here, the threshold is chosen as $\lambda_{n}=\sqrt{n}/\log(n)$ satisfying $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as $n\rightarrow\infty$; $T_{\alpha}$, $T_{\beta}$ and $P_{JS}$ are given in (2.2) and (2.5), respectively. In Figure 3, we provide an illustrative example of the p-values for AJS and JS methods with $(\alpha,\beta)=(0,0)$ and $n=2000$. The data are generated from the same models of Figure 2 with 5000 repetitions. The distribution of $P_{AJS}$ is observed to be uniform in Figure 3, whereas the histogram indicates a right skewness in the distribution of $P_{JS}$. Moreover, the Figure 3 provides an intuitive explanation for the conservatism of traditional JS method when $\alpha=\beta=0$.
@@
The decision rule of AJS test is that the null hypothesis $H_{0}$ is rejected if $P_{AJS}$ is much smaller than the significance level $\delta$. i.e., we can regard $M$ as a significant mediator when $P_{AJS}<\delta$ in practical application, where $P_{AJS}$ is defined by (2.12). Furthermore, we derive the explicit expressions of size and power in Theorem 1, which is useful for ensuring rationality of the AJS test.
@@
Given that $\lambda_{n}=o(\sqrt{n})$ and as the sample size $n$ approaches infinity, with probability approaching one, under $H_{00}$ the asymptotic size of our proposed AJS test for (2.1) is
@@
Theoretically, we compare the AJS with traditional JS in terms of size. First, under $H_{00}$ the size of traditional JS test for (2.1) is
@@
which explains the conservative phenomenon of traditional JS test. Based on (2.13), $Size(AJS|H_{00})$ is much larger than $Size(JS|H_{00})$ as $n\rightarrow\infty$. As a whole, the proposed AJS method is obviously better than the conventional JS test in many situations. Simulations will be used to support this conclusion.
In the field of mediation analysis, another popular method for (2.1) is the Sobel test Sobel (1982), and the corresponding test statistic is
@@
where $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively. $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively. The decision rule of traditional Sobel test relies on the standard asymptotic normality of $T_{Sobel}$. i.e., we can reject $H_{0}$ if the $p$-value, $P_{Sobel}$, is smaller than a specified significance level $\delta$, where
@@
and $\Phi_{N(0,1)}(\cdot)$ is the cumulative distribution function of $N(0,1)$. Due to the convenience of Sobel test in practical application, it has been widely adopted by many research fields, such as the social and biomedical sciences. However, the Sobel test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). From Liu et al. (2022), the Sobel statistic $T_{sobel}$ has an asymptotic normal distribution $N(0,1)$ under $H_{01}$ and $H_{10}$, while its asymptotic distribution is $N(0,1/4)$ in the case of $H_{00}$. Therefore, the performance of traditional Sobel test is conservative when $\alpha=0$ and $\beta=0$. In fact, this is a long-existed and unresolved problem for the Sobel test. Similar to the idea of AJS method, we propose a novel adjusted Sobel (ASobel) test procedure for (2.1), and the p-value is defined as
@@
where $\Phi_{N(0,1/4)}(\cdot)$ is the cumulative distribution function of $N(0,1/4)$; $T_{\alpha}$, $T_{\beta}$ and $\lambda_{n}$ are the same as that of $P_{AJS}$ in (2.12). The decision rule of the ASobel test states that the null hypothesis $H_{0}$ is rejected if the test statistic ($P_{ASobel}$) is significantly smaller than the predetermined significance level $\delta$. The p-values for ASobel and Sobel methods with $(\alpha,\beta)=(0,0)$ and $n=2000$ are illustrated in Figure S.1. The sample data are the same as depicted in Figure 3. The distribution of $P_{ASobel}$ appears to be uniformly distributed in Figure S.1, while the histogram indicates a right-skewed distribution for $P_{Sobel}$. The Figure S.1 further elucidates the inherent conservatism of the traditional Sobel method when both $\alpha$ and $\beta$ are set to zero.
@@
The proof details of (2.19) of ASobel test are presented in the supplemental material. The ASobel test, although superior to the conventional Sobel test, demonstrates inferior performance compared to the AJS method in numerical simulations. Consequently, our focus on hypothesis testing does not heavily emphasize the ASobel method. However, the ASobel plays a crucial role in constructing efficient confidence intervals for mediation effects, which will be presented in Section 4.

### v8

@@ Abstract
Mediation analysis is an important statistical tool in many research fields, where the joint significance test is widely utilized for examining mediation effects. Nevertheless, the limitation of this mediation testing method stems from its conservative Type I error, which reduces its statistical power and imposes certain constraints on its utility. The proposed solution to address this gap is the adjusted joint significance test for one mediator, which introduces a novel data-adjusted approach for assessing mediation effects that showcases significant advancements. The method is specifically designed to be user-friendly, thereby eliminating the necessity for intricate procedures. We further extend the adjusted joint significance test for small-scale mediation hypotheses with family-wise error rate (FWER) control. Additionally, a novel adjusted Sobel-type confidence interval is proposed for the mediation effects, demonstrating significant advancements over conventional Sobel’s method. The effectiveness of our mediation testing and confidence interval estimation is assessed through extensive simulations, and compared against a multitude of existing approaches. Finally, we present the application of the method to three substantive datasets with continuous, binary and time-to-event outcomes, respectively.
Keywords: Adjusted Sobel’s test; Confidence intervals; Multiple mediators; Small-scale mediation hypotheses.
@@ 1 Introduction
Mediation analysis plays an important role in understanding the causal mechanism that an independent variable $X$ affects a dependent variable $Y$ through an intermediate variable (mediator) $M$. The utilization of mediation analysis is extensively prevalent across various disciplines, such as psychology, economics, epidemiology, medicine, sociology, behavioral science, and many others. From the perspective of methodological development, Baron and Kenny (1986) have laid a solid foundation for mediation analysis. Subsequently, numerous studies have been conducted on this subject. Just to name a few: Mackinnon et al. (2004) constructed the confidence limits for indirect effect by resampling methods. Wang and Zhang (2011) and Zhang and Wang (2013) introduced the estimating and testing methods for mediation effects with censored data and missing data, respectively. Shen et al. (2014) proposed an inference technique for quantile mediation effects. VanderWeele and Tchetgen (2017) considered causal mediation analysis with time-varying exposures and mediators. Sun et al. (2021) proposed a Bayesian modeling approach for mediation analysis. Zhou (2022) introduced a semiparametric estimation method for mediation analysis with multiple causally ordered mediators. Zhang and Li (2023) used the subsampled double bootstrap and divide-and-conquer algorithms to conduct statistical mediation analysis on large-scale datasets. He et al. (2024) developed an adaptive bootstrap framework that can be applied to the joint significance test of mediation effect. For more results about mediation analysis, we refer to the reviewing papers by MacKinnon et al. (2007) and Preacher (2015).
@@
The joint significance test is a crucial statistical approach in the field of mediation analysis, which plays a pivotal role in investigating the causal mechanism underlying mediation effects MacKinnon (2008). However, the main shortcoming of this method is due to the conservative type I error of mediation testing MacKinnon et al. (2002), which largely prevents its popularity for practical users. The statistical analysis of the joint significance tests for large-scale mediation hypotheses in genome-wide epigenetic research has been extensively investigated (Huang, 2019; Dai et al., 2022; Liu et al., 2022), whereas these methods are not applicable to single or small-scale mediation hypotheses. Furthermore, the confidence interval for the mediation effect is a crucial aspect in mediation analysis, which is highly valuable in comprehending the mediation mechanism. The literature on mediation analysis lacks a substantial number of studies focusing on confidence interval estimation. The Sobel-type (or normality-based) method and Bootstrap are consistently employed to construct confidence intervals of mediation effects. However, these two methods are inadequate when both the pathway effects along $X\rightarrow M$ and $M\rightarrow Y$ equal zero. To improve the performances of joint significance test and Sobel-type confidence interval, we propose two novel data-adjusted mediation analysis methods with theoretical verification. The main advantages of our proposed method are as follows: First, the adjusted joint significance test and adjusted sobel-type confidence interval are two flexible and data-driven methods. The implementation of the two proposed methods is particularly convenient from a practical perspective. Specifically, our method ensures user-friendliness by eliminating complex procedures. Second, the test method we propose exhibits significant advancements in terms of size and power when compared to the conventional joint significance test. The enhanced powers are particularly evident for those mediation effects that are relatively weak, making them challenging to be recognized as significant mediators by traditional methods. Third, the explicit formulation of the coverage probability and length of the adjusted Sobel-type confidence interval is provided for comparison with conventional Sobel’s method.
@@
The remainder of this paper is organized as follows: In Section 2, we review some details about the traditional joint significance test for mediation effects. Then we propose a novel data-adjusted joint significance test for one mediator, together with the explicit expression of size. Meanwhile, an adjusted Sobel test is also introduced, which shows a significant improvement compared to traditional Sobel’s method. In Section 3, we implement the adjusted joint significance test towards small-scale multiple testing with FWER control. Section 4 introduces an adjusted Sobel-type method to construct confidence interval for the mediation effect. Section 5 presents some simulation studies to assess the performance of our method. In Section 6, we perform mediation analysis for three real-world datasets with the proposed method. Some concluding remarks are provided in Section 7. All proof details of theorems are presented in the Supplementary Material. Finally, our method offers a publicly available and user-friendly R package, called $\tt AdjMed$, which can be accessed at https://github.com/zhxmath/AdjMed.
@@ 2 Adjusted Joint Significance Test
To begin with, we review some fundamental notations within the context of mediation analysis. Let $X$ be an exposure, $M$ be the mediator and $Y$ be the outcome (see Figure 1). As described by MacKinnon (2008), the aim of mediation analysis is focused on investigating the causal mechanism along the pathway $X\rightarrow M\rightarrow Y$. Generally speaking, the causal effect $X\rightarrow M$ is parameterized by $\alpha$ (after adjusting for confounders), and the causal effect $M\rightarrow Y$ is parameterized by $\beta$ (after adjusting for exposure and confounders). The mediating effect of $M$ in this case is described by $\alpha\beta$, commonly known as the “product-coefficient” approach MacKinnon et al. (2002). To evaluate whether $M$ plays an intermediary role between $X$ and $Y$, it is customary to conduct hypothesis testing at a significance level of $\delta$ as follows:
@@
The rejection of the null hypothesis would indicate that $M$ is a statistically significant mediator in the pathway $X\rightarrow M\rightarrow Y$. The null hypothesis $H_{0}$ is worth noting as it is composite. Specifically, $H_{0}$ can be decomposed equivalently into the union of three disjoint component null hypotheses $H_{0}=H_{00}\cup H_{01}\cup H_{10}$, where
@@
Let $T_{\alpha}={\hat{\alpha}}/{\hat{\sigma}_{\alpha}}$ and $T_{\beta}={\hat{\beta}}/{\hat{\sigma}_{\beta}}$ be the statistics for testing $\alpha=0$ and $\beta=0$. Here $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively; $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively. Under the null hypothesis, as the sample size $n$ tends to infinity, it can be observed that
$T_{\alpha}\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}N(0,1)~{}{\rm and}~{}T_{\beta}\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}N(0,1),$ (2.2)
where $\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}$ denotes convergence in distribution. The corresponding p-values for $T_{\alpha}$ and $T_{\beta}$ are
@@
where $T_{\alpha}$ and $T_{\beta}$ are defined in (2.2), $\Phi_{N(0,1)}(\cdot)$ is the cumulative distribution function of $N(0,1)$. The joint significance (JS) test, also known as the MaxP test, is widely recognized as one of the most popular methods for mediation analysis in the field. The purpose of the JS test is to reject $H_{0}$ when both $\alpha=0$ and $\beta=0$ are simultaneously rejected. The JS test statistic is
@@
where $P_{\alpha}$ and $P_{\beta}$ are given in (2.3) and (2.4), respectively. The practical applicability of JS test has led to its widespread adoption across various research fields, including the social and biomedical sciences. However, the JS test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). In the literature, this is a long-existed and unresolved problem for the JS test. To bridge this gap, we aim to propose a novel data-adjusted method for improving the statistical efficiency of the JS test, especially focusing on the conservative issue in the case of $\alpha=\beta=0$.
Under $H_{0}$, the traditional JS test regards $P_{JS}$ as a uniform random variable over $(0,1)$. i.e., $P_{JS}\sim U(0,1)$. However, the actual distribution of $P_{JS}$ is not $U(0,1)$ under the component null hypotheses $H_{00}$, which is the reason of the conservative performance of traditional JS test. To provide further insights on this matter, let us consider $Z_{1}$ and $Z_{2}$ as two independent random variables that follow a uniform distribution $U(0,1)$. Let ${Z}=\max({Z_{1}},{Z_{2}})$ and its density function is $f_{{Z}}(z)$. We can derive the distribution function of $Z$ as
@@
Under $H_{00}$, the squared maximum of two p-values ($P^{2}_{JS}$) has an uniform distribution over (0, 1). This discovery provides insight into resolving the conservative challenge of conventional JS test.
@@
To address the conservatism of JS test, we attempt to differentiate $H_{00}$ from $H_{01}$ and $H_{10}$ reasonably using a specific criterion. Under $H_{00}$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{00})=1$, where $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as the sample size $n\rightarrow\infty$. However, $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{10})=0$ and $\lim_{n\rightarrow\infty}\mathbb{P}(|T_{\alpha}|<\lambda_{n},|T_{\beta}|<\lambda_{n}|H_{01})=0$. The null hypothesis $H_{0}$ is more likely to be regarded as $H_{00}$ instead of $H_{01}$ or $H_{10}$ when $\max\{|T_{\alpha}|,|T_{\beta}|\}<\lambda_{n}$. The approach employed by He et al. (2024) in constructing bootstrap-based testing statistics for evaluating mediation effects should be noted, while our focus lies on the development of innovative testing statistics that eliminate the need for re-sampling procedures. The above-mentioned concept is exemplified through a numerical illustration provided by us. Specifically, we generate a series of random samples from the linear mediation models: $Y=0.5+0.5X+\beta M+\epsilon$ and $M=0.5+\alpha X+e$, where $X$, $\epsilon$ and $e$ follow from $N(0,1)$. The resulting $T_{\alpha}$ and $T_{\beta}$ are obtained by the ordinary least square method. In Figure 2, we present the scatter plots of $T_{max}=\max\{|T_{\alpha}|,|T_{\beta}|\}$ and $\lambda_{n}=\sqrt{n}/\log(n)$ with 1000 repetitions, where the sample size is $n=1000$, the parameters $(\alpha,\beta)$ are chosen as $(\alpha,\beta)=(0,0)$, and (0.25, 0), respectively. The observation from Figure 2 suggests that, under the assumption of $\alpha=\beta=0$, there is a high probability that $T_{max}$ is significantly smaller than $\lambda_{n}$ as $n\rightarrow\infty$. However, $T_{max}$ is asymptotically much larger than $\lambda_{n}$ when $\alpha\neq 0$, and the same conclusion also applies to $\beta\neq 0$.
Motivated by the aforementioned findings, we propose a novel adjusted joint significance (AJS) test procedure for (2.1), where the p-value is defined as
@@
Here, the threshold is chosen as $\lambda_{n}=\sqrt{n}/\log(n)$ satisfying $\lambda_{n}=o(\sqrt{n})$ and $\lambda_{n}\rightarrow\infty$ as $n\rightarrow\infty$; $T_{\alpha}$, $T_{\beta}$ and $P_{JS}$ are given in (2.2) and (2.5), respectively. In Figure 3, we provide an illustrative example of the p-values for AJS and JS methods with $(\alpha,\beta)=(0,0)$ and $n=2000$. The data are generated from the same models of Figure 2 with 5000 repetitions. The distribution of $P_{AJS}$ is observed to be uniform in Figure 3, whereas the histogram indicates a right skewness in the distribution of $P_{JS}$. Meanwhile, the Figure 3 provides an intuitive explanation for the conservatism of traditional JS method when $\alpha=\beta=0$.
@@
The decision rule of AJS test is that the null hypothesis $H_{0}$ is rejected if $P_{AJS}$ is much smaller than the significance level $\delta$, where $P_{AJS}$ is defined in (2.12). The explicit expression of size in Theorem 1 is derived, providing a valuable tool for ensuring the rationality of the AJS test.
@@
Given that $\lambda_{n}=o(\sqrt{n})$ and as the sample size $n$ approaches infinity, with probability approaching one, under $H_{00}$ the asymptotic size of our proposed AJS test is
@@
Theoretically, we conduct a comparative analysis between AJS and traditional JS in terms of size. Under $H_{00}$, the size of traditional JS test for (2.1) is
@@
which explains the conservative phenomenon of traditional JS test.
The Sobel test Sobel (1982) is another widely used method for mediation analysis in the field, and the corresponding test statistic is
@@
where $\hat{\alpha}$ and $\hat{\beta}$ are the estimates for $\alpha$ and $\beta$, respectively; $\hat{\sigma}_{\alpha}$ and $\hat{\sigma}_{\beta}$ are the estimated standard errors of $\hat{\alpha}$ and $\hat{\beta}$, respectively. The decision rule of traditional Sobel test relies on the standard asymptotic normality of $T_{Sobel}$. We can reject $H_{0}$ if the $p$-value, $P_{Sobel}$, is smaller than a specified significance level $\delta$, where
@@
$\Phi_{N(0,1)}(\cdot)$ is the cumulative distribution function of $N(0,1)$, and $T_{Sobel}$ is given in (2.15). The Sobel test suffers from overly conservative type I error, especially when both $\alpha=0$ and $\beta=0$ MacKinnon et al. (2002). From Liu et al. (2022), the Sobel statistic $T_{sobel}$ has an asymptotic normal distribution $N(0,1)$ under $H_{01}$ and $H_{10}$, while its asymptotic distribution is $N(0,1/4)$ in the case of $H_{00}$. Therefore, the performance of traditional Sobel test is conservative when $\alpha=0$ and $\beta=0$. Similar to the idea of AJS method, we propose a novel adjusted Sobel (ASobel) test procedure for (2.1), and the p-value is defined as
@@
where $\Phi_{N(0,1/4)}(\cdot)$ is the cumulative distribution function of $N(0,1/4)$; $T_{Sobel}$ is given in (2.15); $T_{\alpha}$, $T_{\beta}$ and $\lambda_{n}$ are the same as that of $P_{AJS}$ in (2.12). The decision rule of the ASobel test states that the null hypothesis $H_{0}$ is rejected if the test statistic ($P_{ASobel}$) is significantly smaller than the predetermined significance level $\delta$. The p-values for ASobel and Sobel methods with $(\alpha,\beta)=(0,0)$ and $n=2000$ are illustrated in Figure S.1. The sample data are the same as depicted in Figure 3. The distribution of $P_{ASobel}$ appears to be uniformly distributed in Figure S.1, while the histogram indicates a right-skewed distribution for $P_{Sobel}$. The Figure S.1 further elucidates the inherent conservatism of the traditional Sobel method when both $\alpha$ and $\beta$ are set to zero.
@@
The proof details of (2.19) of ASobel test are presented in the Supplemental Material. The ASobel test, although superior to the conventional Sobel test, demonstrates inferior performance compared to the AJS method in numerical simulations. Consequently, our focus on hypothesis testing does not heavily emphasize the ASobel method. However, the ASobel plays a crucial role in constructing efficient confidence intervals for mediation effects, which will be presented in the Section 4.